jhonyy9 (jhonyy9):

- prove that for every n natural numbers grater or equal 2 exist one p and k numbers prim grater or equal 2 such that the below equation is true : (p-1) (k-1) ----- + ----- +1 = n 2 2

jhonyy9 (jhonyy9):

for example : (2-1) (2-1) ----- + ------ +1 = 2 2 2

jhonyy9 (jhonyy9):

(3-1) (3-1) ----- + ----- +1 =3 2 2

jhonyy9 (jhonyy9):

(5-1) (3-1) ----- + ----- +1 = 2 +1 +1 =4 2 2

jhonyy9 (jhonyy9):

(5-1) (5-1) ----- + ----- +1 = 2+2+1=5 2 2

jhonyy9 (jhonyy9):

can be this proven that is true for every natural numbers ?